OpenStudy (anonymous):

Solve the compound inequality. 6x<(with the line) -18 or 9x > (with line ) 81

OpenStudy (anonymous):

\[x \le-3 \] or \[x \ge9\] \[9\le x \le-3\] Use this as a reference only as i often get confused between 'and' and 'or'

OpenStudy (anonymous):

blahh I don't get it

OpenStudy (anonymous):

\[6x \le-18\] \[x \le-3\] (this for the first one, simplify it till you have 1 x by itself) and then do the same to the other one

OpenStudy (anonymous):

\[9x \ge81\]\[x \ge 9\]

OpenStudy (anonymous):

I still don't lol

OpenStudy (aravindg):

Why not?

OpenStudy (aravindg):

Which step do you have doubt?

OpenStudy (anonymous):

idk